Solve for r and interpret the result

X 2,4,6,7,10

Y 8,10,12,6,16


1
Expert's answer
2022-06-13T12:54:35-0400

In order to compute the regression coefficients, the following table needs to be used:


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
 & X & Y & XY & X^2 & Y^2 \\ \hline
 & 2 & 8 & 16 & 4 & 64 \\
 \hdashline
 & 4 & 10 & 40 & 16 & 100 \\
 \hdashline
 & 6 & 12 & 72 & 36 & 144 \\
 \hdashline
 & 7 & 6 & 42 & 49 & 36 \\
 \hdashline
 & 10 & 16 & 160 & 100 & 256 \\
 \hdashline
Sum= & 29 & 52 & 330 & 205 & 600 \\
 \hdashline
\end{array}"




"\bar{X}=\dfrac{1}{n}\sum _{i}X_i=\dfrac{29}{5}=5.8"




"\bar{Y}=\dfrac{1}{n}\sum _{i}Y_i=\dfrac{52}{5}=10.4"




"SS_{XX}=\sum_iX_i^2-\dfrac{1}{n}(\sum _{i}X_i)^2""=205-\dfrac{29^2}{5}=36.8"





"SS_{YY}=\sum_iY_i^2-\dfrac{1}{n}(\sum _{i}Y_i)^2""=600-\dfrac{(52)^2}{5}=59.2"




"SS_{XY}=\sum_iX_iY_i-\dfrac{1}{n}(\sum _{i}X_i)(\sum _{i}Y_i)""=330-\dfrac{29(52)}{5}=28.4"





"r=\dfrac{SS_{XY}}{\sqrt{SS_{XX}SS_{YY}}}=\dfrac{28.4}{\sqrt{36.8(59.2)}}"




"=0.608462"




"0.4<r<0.7"

Moderate positive correlation


"m=slope=\dfrac{SS_{XY}}{SS_{XX}}=\dfrac{28.4}{36.8}=0.7717""n=\bar{Y}-m\bar{X}=10.4-\dfrac{28.4}{36.8}(5.8)=5.9239"



The regression equation is:


"y=5.9239+0.7717x"